# Learning General Representation of 12-Lead Electrocardiogram With a Joint-Embedding Predictive Architecture

### v1

@@ Abstract
We propose a self-supervised learning method for 12-lead Electrocardiogram (ECG) analysis, named ECG Joint Embedding Predictive Architecture (ECG-JEPA). ECG-JEPA employs a masking strategy to learn semantic representations of ECG data. Unlike existing methods, ECG-JEPA predicts at the hidden representation level rather than reconstructing raw data. This approach offers several advantages in the ECG domain: (1) it avoids producing unnecessary details, such as noise, which is common in standard ECG; and (2) it addresses the limitations of naïve L2 loss between raw signals. Another key contribution is the introduction of a special masked attention tailored for 12-lead ECG data, Cross-Pattern Attention (CroPA). CroPA enables the model to effectively capture inter-patch relationships. Additionally, ECG-JEPA is highly scalable, allowing efficient training on large datasets. Our code is openly available https://github.com/sehunfromdaegu/ECG_JEPA.
@@ 1 Introduction
Electrocardiograms (ECG) provide a non-invasive method to measure the electrical activity of the heart over time, serving as a crucial tool for diagnosing various cardiac conditions. While numerous supervised learning models have been developed to detect heart diseases using ECG data [1, 2, 3], these models often face significant performance degradation when applied to data distributions different from those on which they were trained.
Self-supervised learning (SSL) has emerged as a powerful paradigm for learning general representations across various domains, including natural language processing [4, 5, 6], computer vision [7, 8, 9], and video analysis [10, 11]. Despite its promise, applying SSL to ECG data presents unique challenges. For instance, data augmentation, which is essential in many SSL architectures, is more complex for ECG than for computer vision (CV) data. Simple transformations like rotation, scaling, and flipping, effective in CV, can distort the physiological meaning of ECG signals. Additionally, ECG recordings often contain artifacts and noise, which cause autoencoder-based SSL models to struggle with reconstructing raw signals. These architectures may also miss visually subtle but diagnostically critical features, such as P-waves and T-waves, which are imperative for diagnosing certain cardiac conditions.
In this work, we propose a novel ECG Joint-Embedding Predictive Architecture (ECG-JEPA) tailored for 12-lead ECG data, effectively addressing the aforementioned challenges. ECG-JEPA utilizes a transformer architecture to capture the semantic meanings of ECG. By masking several patches of the ECG data, ECG-JEPA predicts the abstract representations of the missing patches, indicating a high-level understanding of the data. Additionally, we develop a novel masked-attention for multi-lead ECG data, coined Cross-Pattern Attention (CroPA). CroPA enables the transformer architecture to effectively capture the relationships between patches in multi-lead ECG.
In this work, we demonstrate that:
  • ECG-JEPA significantly improves linear evaluation and fine-tuning on classification tasks compared to existing SSL methods without hand-crafted augmentations.
  • CroPA enhances the model’s ability to capture inter-lead relationships, leading to better performance on downstream tasks.
@@ 2 Background
Self-Supervised Learning (SSL) facilitates learning abstract representations from input data without the need for labeled data, which is particularly beneficial in medical domains where labeled data is scarce and expensive. SSL leverages inherent data patterns to learn useful representations, allowing models to adapt to various downstream tasks with greater robustness to data imbalances [12].
@@ 2.3 Electrocardiogram (ECG)
In this work, we use only 8 leads (I, II, V1-V6) as the remaining 4 leads (III, aVR, aVL, aVF) can be derived from linear combinations of the 8 leads. This choice maintains the necessary diagnostic information while optimizing computational efficiency.
@@ 3 Methodology
ECG-JEPA is trained by predicting masked representations of ECG data within the hidden representation space, using only a partial view of the input. The proposed architecture utilizes a student-teacher network framework, as illustrated in Figure 3. We subdivide the multi-channel ECG into non-overlapping patches and sample a subset of these patches for masking. However, reconstructing the raw signals of masked patches can be particularly challenging in the ECG domain due to the prevalence of noise in biological signals. Instead, our model predicts the masked patches in the hidden representation space, where this challenge can be effectively addressed. We validate the quality of the learned representations through various downstream tasks, including linear probing, fine-tuning on classification tasks, and ECG feature extraction tasks.
@@ 3.1 Patch Masking
The patches are then converted into sequences of token vectors using a linear layer, and augmented with positional embeddings. For positional encoding, we employ the conventional 2-dimensional sinusoidal positional embeddings for the student and teacher networks, while we use 1-dimensional sinusoidal positional embeddings for the predictor network.
Figure 3: ECG-JEPA. The student network processes only non-masked patches of time series, whereas the teacher network processes the entire time series. Then the predictor predicts the representation of masked patches processed by the teacher. Note that both student and teacher network process the multi-channel time series, while predictor gets only a single-channel time series. In this picture, there are $L=3$ leads with $N=5$ subintervals and $Q=3$ unmasked subintervals.
@@ 3.2 Teacher, Student, and Predictor
ECG-JEPA consists of three main components: the teacher network, the student network, and the predictor network. Both the teacher and student networks are based on standard transformer architectures. The weights of the teacher network are updated through an exponential moving average (EMA) of the student network, with details provided in the supplementary materials. The predictor network, a smaller transformer, operates on single-channel representations.
@@ 3.3 Cross-Pattern Attention (CroPA)
Biological signals often require simultaneous observation across multiple channels. In a 12-lead ECG, abnormalities must be consistent across multiple leads to diagnose specific cardiac pathologies. Therefore, the meaning of a single patch should be understood in the context of other patches within the same channel and across different channels. In contrast, patches in different channels and temporal spaces are often neglected when interpreting a 12-lead ECG.
To address this, we introduce Cross-Pattern Attention (CroPA), a masked self-attention tailored for multi-lead ECG data. CroPA allows each patch to attend only to patches within the same channel and temporal space (Figure 4). This method enhances our model’s effectiveness compared to the standard attention mechanism, where patches attend to all other patches indiscriminately.
@@ 4 Related Work
This section provides an overview of existing SSL methods that have been applied to ECG data. We compare ECG-JEPA with these methods on various downstream tasks in the next section.
@@ 4.2 Generative Architectures
BERT [4] pioneered generative architectures in natural language processing (NLP), inspiring similar methods in other domains. Masked Autoencoders (MAE) [8] have been developed for computer vision, where a portion of the input image is masked, and the model is trained to reconstruct these missing parts. Recently, [17] pointed out that generative architectures often prioritize learning the principal subspaces of the data, which may limit their effectiveness in capturing semantic representations for perceptual tasks.
@@ 4.3 Joint-Embedding Predictive Architectures
Data2vec [18] introduced an SSL architecture that predicts the average of the encoder’s final layers’ outputs, which was further improved in efficiency in data2vec 2.0 [19]. Notable models like I-JEPA [9] and V-JEPA [11] predict the hidden representations of masked patches, enabling semantic representation learning for images and videos, respectively.
@@ 4.4 Self-Supervised Learning for ECG
Several attempts have been made to train SSL models for 12-lead ECG data. Contrastive Multi-segment Coding (CMSC) [20] is a contrastive method that divides an ECG into two segments, encouraging close representations for compatible segments while pushing apart incompatible ones. Contrastive Predictive Coding (CPC) [15] has also been used for ECG data in [21], predicting future representations of the ECG. A significant disadvantage of CPC is that it is not scalable to large datasets, as the LSTM modules inside CPC make the training process extremely slow. Recently, [22] utilized masked autoencoders for ECG data, proposing temporal and channel masking strategies, coined as Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE), respectively. Similarly, [23] proposed ST-MEM, an MAE for ECG data that masks random time intervals for each lead. One possible drawback of MLAE and ST-MEM is that predicting some patches can be relatively easy considering that 12 leads signals are highly correlated.
@@ 5 Experimental Settings
This section provides an overview of the experimental settings. Details of all experiments can be found in the supplementary materials.
@@ 5.1 Pretraining Datasets
Training SSL models with large datasets is crucial for developing generalized representations. However, most previous works have used relatively small datasets, with the exception of [23], where an SSL model was trained with a large number of 12-lead ECGs. Following [23], we use the Chapman [24], Ningbo [25], and CODE-15 [26] datasets for pretraining ECG-JEPA. The Chapman and Ningbo datasets collectively consist of 45,152 10-second 12-lead ECGs at 500Hz. CODE-15 includes 345,779 12-lead ECGs from 233,770 patients at 400Hz, with 143,328 being 10-second recordings. After excluding recordings with missing values, we have 43,240 ECGs from Chapman and Ningbo and 130,900 ECGs from CODE-15. All ECGs were resampled to 250Hz for consistency.
@@ 5.2 Downstream Datasets
We use the PTB-XL [27] and CPSC2018 [28] datasets to evaluate the performance of ECG-JEPA on downstream tasks. PTB-XL contains 21,837 clinical 10-second 12-lead ECG records from 18,885 patients, recorded at 500Hz and annotated with 71 diagnostic labels, which are aggregated into five superclasses. We use these superclass labels for our experiments. The CPSC2018 dataset includes 6,877 12-lead ECG recordings with nine annotated cardiac conditions. These datasets are multi-label in nature, where each recording can have multiple labels simultaneously. The details of the datasets are provided in the supplementary materials.
5.3 Pretraining
After resampling the ECG data to 250Hz, each 10-second multi-lead ECG has $T=2500$ time points. We divide the interval $[0,T]$ into $N=50$ non-overlapping subintervals of length $t=50$. The model is trained for 100 epochs without any data augmentations, and the final checkpoint is used for downstream tasks.
@@ 5.4 Architecture
We use vision transformer (ViT-Base) architectures [29] for the student and teacher network, and a smaller standard transformer architecture for the predictor network. The student and teacher network process the multi-lead ECG, while the predictor processes each lead independently to reconstruct the masked representations. See supplementary materials for detailed architecture specifications.
@@ 5.5 Downstream Tasks
We conduct extensive experiments to show that ECG-JEPA effectively captures semantic representations. We evaluate its performance on classification tasks through linear probing and fine-tuning. Additionally, we evaluate the performance on the low-shot setting, and on the reduced-lead setting where the downstream dataset is restricted to have a single or two leads.
Next, we validate the expressive power of the learned representations by predicting ECG features, including heart rate and QRS duration. To the best of our knowledge, this is the first work showing that the learned representations can recover various ECG features. The successful prediction of these features provides compelling evidence that the representations are not only informative but also capture essential and clinically relevant characteristics, making them highly reliable for ECG analysis.
In ECG analysis, it is common for a single recording to have multiple labels simultaneously, making datasets like PTB-XL and CPSC2018 multi-label tasks. However, many previous studies have converted this into a multi-class classification problem by restricting the dataset to subsets with only a single label. To ensure a fair comparison, we will pretrain the methods using publicly available code and evaluate their performance on the multi-label classification task. If the code is not available, we will convert our task into a multi-class problem and compare the results with the reported performance in the literature.
@@ 6 Experiments
In this section, we assess the performance of the learned representations on various downstream tasks. Our experiments aim to demonstrate that the learned representations are generalizable and effective in capturing essential ECG features. We compare ECG-JEPA with several state-of-the-art SSL methods.
For classification tasks, we use AUC (Area Under the ROC Curve) and F1 scores as our primary evaluation metrics. AUC evaluates model performance across all possible thresholds, providing a comprehensive and stable measure of discriminative ability. In contrast, the F1 score balances precision and recall at a fixed threshold. In our experiments, we use the AUC as our main metric, as it provides a more robust evaluation of the model’s performance across varying classification thresholds, ensuring that the results are not overly dependent on a single decision boundary.
For multi-label classification, the AUC is computed by averaging the scores of binary classifications for each label. For multi-class classification, the AUC is calculated using the average of one-vs-rest approach. In both cases, the F1 score is computed using a macro-averaged approach that averages the F1 scores of all classes.
In most cases, ECG-JEPA outperforms other SSL methods that leverage hand-crafted augmentations, demonstrating its effectiveness in learning general representations of ECG data. In experiments, E-JEPArb and E-JEPAmb refer to ECG-JEPA models trained with random masking and multi-block masking strategies, respectively.
6.1 Linear Evaluation
Tables 1 and 2 present the results of our linear evaluation on the PTB-XL and CPSC2018 datasets. We train a linear classifier on top of the frozen representations for 10 epochs, and evaluate its performance on downstream tasks. As shown in the tables, ECG-JEPA consistently outperforms other SSL methods, demonstrating superior efficiency and effectiveness with substantially reduced computational resources.
@@ 6.1 Linear Evaluation
MTAE [^1] Score as reported in [23]. 800 0.807 0.437 0.818 0.349
MLAE [^1] Score as reported in [23]. 800 0.779 0.382 0.794 0.263

### v2

@@ Abstract
We propose a self-supervised learning method for 12-lead Electrocardiogram (ECG) analysis, named ECG Joint Embedding Predictive Architecture (ECG-JEPA). ECG-JEPA employs a masking strategy to learn semantic representations of ECG data. Unlike existing methods, ECG-JEPA predicts at the hidden representation level rather than reconstructing raw data. This approach offers several advantages in the ECG domain: (1) it avoids producing unnecessary details, such as noise, which is common in standard ECG; and (2) it addresses the limitations of naïve L2 loss between raw signals. Another key contribution is the introduction of a special masked attention tailored for 12-lead ECG data, Cross-Pattern Attention (CroPA). CroPA enables the model to effectively capture inter-patch relationships. Additionally, ECG-JEPA is highly scalable, allowing efficient training on large datasets. Our code is openly available at https://github.com/sehunfromdaegu/ECG_JEPA.
@@ 1 Introduction
Electrocardiograms (ECG) provide a non-invasive method to measure the electrical activity of the heart over time, serving as a crucial tool for diagnosing various cardiac conditions. While numerous supervised methods have been developed to detect heart diseases using ECG data [1, 2, 3], these models often face significant performance degradation when applied to data distributions different from those on which they were trained.
Self-supervised learning (SSL) has emerged as a powerful paradigm for learning general representations across various domains, including natural language processing (NLP) [4, 5, 6], computer vision (CV) [7, 8, 9], and video analysis [10, 11]. Despite its promise, applying SSL to ECG data presents unique challenges. For instance, data augmentation, which is essential in many SSL architectures, is more complex for ECG than for computer vision data. Simple transformations like rotation, scaling, and flipping, effective in CV, can distort the physiological meaning of ECG signals. Additionally, ECG recordings often contain artifacts and noise, which cause autoencoder-based SSL models to struggle with reconstructing raw signals. These architectures may also miss visually subtle but diagnostically critical features, such as P-waves and T-waves, which are imperative for diagnosing certain cardiac conditions.
In this work, we propose a novel ECG Joint-Embedding Predictive Architecture (ECG-JEPA) tailored for 12-lead ECG data, effectively addressing the aforementioned challenges. ECG-JEPA utilizes a transformer architecture to capture the semantic meanings of ECG. By masking several patches of the ECG data, ECG-JEPA predicts the abstract representations of the missing patches, indicating a high-level understanding of the data. Additionally, we develop a novel masked-attention for multi-lead ECG data, coined Cross-Pattern Attention (CroPA). CroPA incorporates clinical knowledge into the model as an inductive bias, guiding it to focus on clinically relevant patterns and relationships across leads.
We demonstrate the following contributions of ECG-JEPA:
  • ECG-JEPA achieves notable improvements in linear evaluation and fine-tuning on classification tasks compared to existing SSL methods without hand-crafted augmentations.
  • CroPA introduces a specialized masked attention mechanism, allowing the model to focus on clinically relevant information in multi-lead ECG data, resulting in improved downstream task performance.
@@ 2 Background
Self-Supervised Learning (SSL) facilitates learning abstract representations from input data without the need for labeled data, which is particularly beneficial in medical domains where labeled data is scarce and expensive. SSL leverages inherent data patterns to learn useful representations, allowing models to adapt to various downstream tasks with greater robustness to data imbalances [12]. In this section, we briefly explain key SSL techniques in Sections 2.1 and 2.2, including generative architectures and joint-embedding predictive architectures. Additionally, we provide an overview of the ECG and its key features in Sections 2.3 and 2.4, highlighting the critical characteristics essential for understanding ECG data.
@@ 2.3 Electrocardiogram (ECG)
In this work, we use only 8 leads (I, II, V1-V6) as the remaining 4 leads (III, aVR, aVL, aVF) can be derived from linear combinations of the 8 leads (see C). This choice maintains the necessary diagnostic information while optimizing computational efficiency.
@@ 2.4 ECG Features
Figure 3: ECG-JEPA. The teacher network processes the complete ECG data, generating fully contextualized representations. In contrast, the student network receives only the non-masked patches. The student’s representations are then passed to the predictor, which reconstructs the fully contextualized representations produced by the teacher. The predicted 6 representations are then compared with the corresponding teacher’s representation. This figure illustrates $L=3$ leads with $N=5$ subintervals and $Q=3$ unmasked subintervals.
@@ 3 Methodology
ECG-JEPA is trained by predicting masked representations of ECG data in the hidden representation space, using only a partial view of the input. The proposed architecture utilizes a student-teacher framework, as illustrated in Figure 3. We subdivide the multi-channel ECG into non-overlapping patches and sample a subset of these patches for masking. However, reconstructing the raw signals of masked patches can be particularly challenging in the ECG domain due to the prevalence of noise in biological signals. Instead, our model predicts the masked patches in the hidden representation space, where this challenge can be effectively addressed. We validate the quality of the learned representations through various downstream tasks, including linear probing, fine-tuning on classification tasks, and ECG feature extraction tasks.
@@ 3.1 Patch Masking
The patches are converted into sequences of token vectors using a linear layer, and augmented with positional embeddings. We employ the conventional 2-dimensional sinusoidal positional embeddings for the student and teacher networks, while we use 1-dimensional sinusoidal positional embeddings for the predictor network.
@@ 3.2 Teacher, Student, and Predictor
ECG-JEPA consists of three main components: the teacher network, the student network, and the predictor network. Both the teacher and student networks are based on standard transformer architectures. The weights of the teacher network are updated using an exponential moving average (EMA) of the student network, as detailed in B. The predictor network, a smaller transformer, operates on single-channel representations, which still encode information from all leads due to the self-attention mechanism.
@@ 3.3 Cross-Pattern Attention (CroPA)
Multi-lead ECG signals require careful analysis of patterns that are often consistent across different leads, which is crucial for identifying potential cardiac abnormalities. This demands attention mechanisms that prioritize relationships within the same lead and within relevant time windows.
To incorporate this structural insight, we introduce Cross-Pattern Attention (CroPA), a masked self-attention mechanism designed for multi-lead ECG data. CroPA imposes an inductive bias by allowing each patch to attend only to patches within the same lead and temporal space (Figure 4). This aligns with the way ECG signals are typically interpreted, where intra-lead and temporally adjacent signals hold the most significance.
By incorporating this inductive bias, CroPA helps the model focus on relevant intra-lead relationships, reducing interference from unrelated signals across different channels and time points. Compared to the standard self-attention mechanism, which treat all patches equally, CroPA reflects a structured approach that mirrors the process of multi-lead signal interpretation, leading to improved performance in downstream tasks.
@@ 4 Related Work
This section provides an overview of existing SSL methods that have been applied to ECG data. We compare ECG-JEPA with these methods on various downstream tasks in Section 6.
@@ 4.2 Generative Architectures
Generative architectures, first popularized by BERT [4] in natural language processing (NLP), have since inspired a variety of approaches in other domains. For example, in computer vision, Masked Autoencoders (MAE) [8] extend this concept by masking portions of an input image and training the model to reconstruct the missing parts, similar to how BERT predicts masked tokens in text. Recently, the authors of [17] pointed out that generative architectures often prioritize learning the principal subspaces of the data, which may limit their effectiveness in capturing semantic representations for perceptual tasks.
@@ 4.3 Joint-Embedding Predictive Architectures
In contrast to generative approaches, which focus on reconstructing data in the input space, joint-embedding predictive architectures (JEPA) aim to predict in the representation space, allowing models to learn more abstract and semantic features. An example of this approach is Data2vec [18], which introduced an SSL architecture that predicts the average of the encoder’s final layer outputs. Other examples include I-JEPA [9] and V-JEPA [11], which predict the hidden representations of masked patches, enabling semantic representation learning for images and videos, respectively.
@@ 4.4 Self-Supervised Learning for ECG
Various SSL models have been explored for 12-lead ECG data. Contrastive Multi-segment Coding (CMSC) [19] splits an ECG into two segments, encouraging similar representations for compatible segments while separating incompatible ones. Contrastive Predictive Coding (CPC) [15], applied in [20], predicts future ECG representations, but its reliance on LSTM modules makes it inefficient for large datasets. More recently, [21] introduced masked autoencoders for ECG, proposing temporal and channel masking strategies, Masked Time Autoencoder (MTAE) and Masked Lead Autoencoder (MLAE). Similarly, [22] proposed ST-MEM, which masks random time intervals for each lead. However, both MLAE and ST-MEM may struggle with the high correlations between ECG leads, making some predictions too easy.
@@ 5 Experimental Settings
In all experiments, 10-second multi-lead ECG signals were resampled to 250Hz, yielding $T=2500$ time points. We divided the interval $[0,T]$ into $N=50$ non-overlapping subintervals, each of length $t=50$. The model was trained for 100 epochs without data augmentation, and the final checkpoint was used for downstream tasks. Additional experimental details are provided in Appendix A.
@@ 5.1 Pretraining Datasets
Training SSL models with large datasets is crucial for developing generalized representations. However, most previous works have used relatively small datasets, with the exception of [22], where an SSL model was trained with a large number of 12-lead ECGs. Following [22], we use the Chapman [23], Ningbo [24], and CODE-15 [25] datasets for pretraining ECG-JEPA. The Chapman and Ningbo datasets collectively consist of 45,152 10-second 12-lead ECGs at 500Hz. CODE-15 includes 345,779 12-lead ECGs from 233,770 patients at 400Hz, with 143,328 being 10-second recordings. After excluding recordings with missing values, we have 43,240 ECGs from Chapman and Ningbo and 130,900 ECGs from CODE-15.
@@ 5.2 Downstream Datasets
We use the PTB-XL [26] and CPSC2018 [27] datasets to evaluate the performance of ECG-JEPA on downstream tasks. PTB-XL contains 21,837 clinical 10-second 12-lead ECG records from 18,885 patients, recorded at 500Hz and annotated with 71 diagnostic labels, which are aggregated into five superclasses. We use these superclass labels for our experiments. The CPSC2018 dataset includes 6,877 12-lead ECG recordings with nine annotated cardiac conditions. These datasets are multi-label in nature, where each recording can have multiple labels simultaneously. The details of the datasets are provided in Appendix A.3.
@@ 5.3 Architecture
Our model employs transformer encoder architectures for the student, teacher, and predictor networks. Both the teacher and student networks consist of 12 layers with 16 attention heads and a hidden dimension of 768. The predictor network, designed as a smaller transformer encoder, comprises 6 layers with 12 attention heads and a hidden dimension of 384. While the teacher and student networks process the multi-lead ECG data holistically, the predictor operates on each lead independently to reconstruct the masked representations. Importantly, this does not imply that the predictor relies solely on single-lead information for the reconstruction task; due to the self-attention mechanism, the input representations for each lead still encapsulate information from all leads.
@@ 5.4 Downstream Tasks
We conduct extensive experiments to show that ECG-JEPA effectively captures semantic representations. Its performance is evaluated on classification tasks using linear probing and fine-tuning. Furthermore, we assess its capability in low-shot learning settings, as well as under reduced-lead conditions where the downstream dataset is limited to single or two leads. Reduced-lead configurations are common in clinical practice, especially in scenarios like wearable devices or remote monitoring, where using the full 12-lead ECG setup is impractical.
To validate the expressiveness of the learned representations, we predict key ECG features such as heart rate and QRS duration. Notably, this work is the first to show that these learned representations can recover a variety of ECG features. The ability to predict these features highlights the informativeness of the representations and their potential to capture clinically relevant characteristics, which is crucial for reliable ECG analysis.
ECG datasets, such as PTB-XL and CPSC2018, often include multiple simultaneous labels for a single recording, making them multi-label tasks. However, many prior studies have simplified this into a multi-class classification problem by focusing on single-label subsets of the data. To ensure a fair comparison, we pretrain competing methods using publicly available code and evaluate them on the multi-label classification task. In cases where the code is unavailable, we will convert our task into a multi-class problem to align with the reported performance in the literature.
@@ 6 Experiments
In this section, we evaluate the performance of the learned representations across various downstream tasks to demonstrate their generalizability and ability to capture essential ECG features. ECG-JEPA is compared against several state-of-the-art self-supervised learning (SSL) methods.
For classification tasks, we use AUC (Area Under the ROC Curve) and F1 scores as evaluation metrics. AUC provides a comprehensive measure of discriminative ability by considering performance across all classification thresholds, making it more robust to variations in decision boundaries. In contrast, the F1 score balances precision and recall at a fixed threshold, offering insights into the model’s performance when a specific decision boundary is chosen.
In multi-label classification, we compute AUC by averaging the scores from binary classification for each label, while for multi-class classification, AUC is calculated using the one-vs-rest approach. For both tasks, F1 scores are macro-averaged across all classes to ensure equal weighting of each class in the final score.
In most cases, ECG-JEPA consistently outperforms other SSL methods that rely on hand-crafted augmentations, highlighting its effectiveness in learning generalizable representations from ECG data. In our experiments, E-JEPArb and E-JEPAmb refer to ECG-JEPA models trained using random masking and multi-block masking strategies, respectively.
@@
MTAE111 800 0.807 0.437 0.818 0.349
MLAE111 800 0.779 0.382 0.794 0.263
@@
6.1 Linear Evaluation
Tables 1 and 2 present the results of our linear evaluation on the PTB-XL and CPSC2018 datasets. We train a linear classifier on top of the frozen representations for 10 epochs and evaluate its performance on downstream tasks. Further training beyond 10 epochs does not lead to any significant improvement in performance. As shown in the tables, ECG-JEPA consistently outperforms other SSL methods, demonstrating superior efficiency and effectiveness with substantially reduced computational resources.